### v3

# WW-FL: Secure and Private Large-Scale Federated Learning

Abstract.
Federated learning (FL) is an efficient approach for large-scale distributed machine learning that promises data privacy by keeping training data on client devices. However, recent research has uncovered vulnerabilities in FL, impacting both security and privacy through poisoning attacks and the potential disclosure of sensitive information in individual model updates as well as the aggregated global model. This paper explores the inadequacies of existing FL protection measures when applied independently, and the challenges of creating effective compositions.
Addressing these issues, we propose WW-FL, an innovative framework that combines secure multi-party computation (MPC) with hierarchical FL to guarantee data and global model privacy. One notable feature of WW-FL is its capability to prevent malicious clients from directly poisoning model parameters, confining them to less destructive data poisoning attacks. We furthermore provide a PyTorch-based FL implementation integrated with Meta’s CrypTen MPC framework to systematically measure the performance and robustness of WW-FL. Our extensive evaluation demonstrates that WW-FL is a promising solution for secure and private large-scale federated learning.
Federated Learning, Secure Multi-party Computation, Privacy, MPC, Hierarchical FL, Poisoning, Attacks and Defenses
1. Introduction
Federated learning (FL) (McMahan et al., 2017) is a revolutionary approach for large-scale distributed machine learning that emerged in the past decade and has had a profound impact on both academia (Yin et al., 2022) and industry (Rieke et al., 2020; Ludwig et al., 2020; He et al., 2020). A World Wide Web Consortium (W3C) community group currently works towards establishing FL-related Web standards (W3C, 2023). FL shares some objectives with privacy-preserving machine learning (PPML) (Gilad-Bachrach et al., 2016; Mohassel and Zhang, 2017; Cabrero-Holgueras and Pastrana, 2021), but differs in its approach to training a model: In PPML, training on the entire set of available data is either outsourced to a remote party or done collaboratively while ensuring data privacy. In contrast, FL involves multiple rounds of training, where each round consists of selected clients locally training a model on their private data and a central party performing aggregation of these models to update a global model.
Initially, FL appeared to indeed offer privacy since the training data remains on the client devices, and only model updates (so-called gradients) are transmitted. However, subsequent research has demonstrated that these gradients still reveal a significant amount of information, enabling the central aggregator to infer sensitive details about the clients’ private training data (Zhu et al., 2019; Geiping et al., 2020). To address this issue, secure aggregation (SA) schemes (Bonawitz et al., 2017; Bell et al., 2020) were proposed, where only the aggregated result is revealed to the central party, effectively concealing the individual gradients. Unfortunately, FL poses additional challenges that require careful attention.
Figure 1. Visual representation of deploying our secure world-wide FL (WW-FL) framework. In this illustration, clients connect to local MPC clusters, which in turn are connected to three global servers (this can be varied as per requirements) strategically positioned across continents. Fig. 2 provides a detailed illustration of WW-FL, showcasing the roles of involved entities.
The first challenge relates to serious privacy vulnerabilities that have been identified when using secure aggregation with a single aggregator (So et al., 2021; Boenisch et al., 2021, 2022; Fowl et al., 2022; Wen et al., 2022): Recent studies like Mandrake (Zhao et al., 2023) have demonstrated the effectiveness of privacy attacks even when utilizing secure aggregation techniques across a large number of clients, showing that such schemes are insufficient (Pasquini et al., 2022; Boenisch et al., 2023). This hinders utilizing standard FL and even FL with SA for global, worldwide training tasks as sharing sensitive data across jurisdictions in many cases is legally problematic. For example, the Court of Justice of the European Union (CJEU) with its famous “Schrems II” judgment has ruled the “Privacy Shield” agreement between the EU and the US for exchanging personal data invalid (Tracol, 2020). Until agreeing on the most recent “Data Privacy Framework” (U.S. Department of Commerce, 2022), numerous attempts to set up data sharing have failed, including the “Safe Harbour” principles (Weiss and Archick, 2016). Recent works therefore suggest a distributed aggregator setup based on secure multi-party computation (MPC) techniques to securely distribute aggregation tasks among multiple servers, ensuring privacy despite potential collusion among a subset of them (Fereidooni et al., 2021; Ben-Itzhak et al., 2022; Rathee et al., 2023; Gehlhar et al., 2023).
The second (and closely related) challenge is the current lack of global model privacy. Initially, FL focused on enhancing user participation and improving model accuracy rather than safeguarding the model itself. However, with the widespread adoption of FL in various industries, such as the medical sector (Rieke et al., 2020), the need for privacy protection becomes apparent. Unfortunately, unrestricted access to the aggregated model allows extraction of traces of original training data, necessitating global model privacy (Pasquini et al., 2022; Boenisch et al., 2023; Zhao et al., 2023). So far, only a small number of works have focused on preserving the privacy of the global model in FL (Mandal and Gong, 2019; Sav et al., 2021). Most of these works use homomorphic encryption (HE) schemes (Acar et al., 2018), which primarily address semi-honest corruptions. Moreover, several works do not consider the collusion of clients with the aggregator and thereby provide a weaker notion of security.
The third challenge arises when corrupted clients intentionally manipulate the locally trained model to reduce the system’s accuracy or to introduce backdoors to extract sensitive information in future rounds (Wang et al., 2020; Xie et al., 2020). This method of model poisoning is more potent than data poisoning (Shejwalkar et al., 2022; Xie et al., 2024), which involves manipulating the underlying dataset. Although numerous defenses have been proposed against model poisoning attacks (Shejwalkar and Houmansadr, 2021; Cao et al., 2021), the most effective ones remain impractical for large-scale deployment due to their significant computation and communication costs (Nguyen et al., 2022). Additionally, recent research has introduced even more potent model poisoning strategies that can circumvent state-of-the-art defenses (Shejwalkar and Houmansadr, 2021; Xie et al., 2024). For instance, PoisonedFL (Xie et al., 2024) breaks eight state-of-the-art defenses without any knowledge of the honest clients’ models, highlighting the urgent need for new defense mechanisms. To address corrupted clients, we limit their attack capabilities to data poisoning rather than the more severe model poisoning. This allows us to implement more moderate defenses compared to existing schemes (cf. §5).
Besides these three challenges, a large-scale, worldwide FL deployment needs to account for client heterogeneity in terms of bandwidth, jurisdiction, trust assumptions, and computational power. To elaborate, trust in government and private entities varies widely, influencing how data is shared and used. For instance, strict data privacy laws in some countries impose heavy fines for non-compliance, while data breaches in companies like TalkTalk [^1] https://www.talktalkgroup.com and Target [^2] https://target.com/ have led to significant losses in customer trust and revenue (Erin, 2022; Hibberd, 2022). In the context of training machine learning models for sensitive applications, such as medical diagnostics, diverse patient data is crucial but must be handled with strict privacy.
Although various works have addressed the aforementioned issues individually or in combination, a comprehensive solution remains elusive. Therefore, we propose a unified framework called WW-FL, which enables secure and private distributed machine learning at scale. Departing from traditional FL, we introduce the concept of a trust zone, allowing data to be shared securely within a trusted legal jurisdiction and leveraging privacy-enhancing technologies like MPC. This approach supports edge nodes with limited resources and encourages wider participation, enhancing the system’s overall privacy and performance. We give a more detailed discussion and additional motivation in App. A.1.
Our Contributions
In this paper, we initiate a discussion on novel approaches towards secure and private large-scale federated learning to address the challenges outlined above. Towards this goal, we propose a unified framework called “worldwide-FL” (WW-FL). Our framework is based on a novel abstraction that also captures existing hierarchical FL architectures in a hybrid manner as visualized in Fig. 1 and detailed in §3.5.
Table 1. Comparison of WW-FL and previous works. Notations: $\mathcal{S}$ – Aggregation Server(s), $\mathcal{C}$ – Client, GM – Global Model, LM – Local Model, MPC – Secure Multi-party Computation, DP – Differential Privacy. Given the extensive literature, this comparison focuses on one representative work per category.
Categories Representative Work Method Privacy ($\mathcal{S}$) Privacy ($\mathcal{C}$) Defense Cross Device No Client Interaction Dropout Handling
GM LM GM
Aggregation (Plain) (McMahan et al., 2017)
Aggregation (Robust) (Yin et al., 2018)
Secure Aggregation (Single $\mathcal{S}$) (Bonawitz et al., 2017) Masking
Secure Aggregation (Multi $\mathcal{S}$) (Nguyen et al., 2022) MPC
Hierarchical FL (HFL) (Yang, 2021) DP
Worldwide FL (WW-FL) This work MPC
Briefly, in our framework, FL clients use secret-sharing to securely outsource training data to distributed cluster servers that are based on MPC. The clients then can leave and only sporadically return to provide more training data – this makes our framework robust against real-world issues such as drop-outs, requires no interaction between clients, and resource-constrained (mobile or edge) devices have very little workload. The models, which are trained by the cluster servers with MPC-based PPML techniques, are then aggregated across all training clusters using one or multiple levels of distributed aggregators. For secure distributed aggregation, we again utilize MPC. Note that after aggregation, models are not publicly released but in secret-shared form handed back to the training clusters for the next training iteration. After training is completed, known secure inference protocols can be used to allow private queries (Mann et al., 2023) in a controlled (potentially rate-limited) way. This architecture design addresses issues with single aggregators by relying on a distributed aggregator and notably achieves global model privacy due to a novel combination of FL with PPML techniques.
We observe that a unique property of our framework is the strictly limited attack surface: similar to PPML, malicious clients are restricted to weaker data-poisoning attacks as there is no possibility to access and manipulate the model itself. As defending against data poisoning in PPML is highly non-trivial, we show that state-of-the-art data-poisoning attacks in the suggested hierarchical configuration are less effective than in plain FL. Furthermore, we consider different robust aggregation schemes to further mitigate the effect of such attacks. For this, we additionally propose new heuristics that improve the efficiency in MPC. Specifically, we propose a lightweight variant of the “Trimmed Mean” (Yin et al., 2018) defense for our setting with only data-poisoning attacks.
Finally, we systematically evaluate possible instantiations of WW-FL in terms of performance (both in terms of achievable accuracy as well as overhead due to MPC) and resilience against poisoning attacks. For this, we implement a prototype based on PyTorch and Meta’s CrypTen framework (Knott et al., 2021) for instantiating MPC. We then evaluate the performance when training neural networks for standard image classification tasks in realistic network settings and using GPU-accelerated AWS EC2 instances. This way, we show that the computation and communication overhead induced by MPC is feasible even for large-scale deployments. For example, one round of training LeNet privately with security against semi-honest servers on the MNIST dataset for 5 epochs with batch size 80 takes less than 5 minutes for a cluster with 10 clients, and global aggregation among 10 clusters takes less than 1 second. Moreover, we find that the impact of MPC-induced inaccuracies is less than 0.1%, and models trained with WW-FL converge significantly faster than in plain FL.
In summary, we provide the following contributions:
  • $\bullet$
    We initiate a discussion to explore why novel approaches are necessary to jointly address concerns regarding security, privacy, and scalability in FL.
  • $\bullet$
    We propose a new generic (hierarchical) FL framework called WW-FL that achieves complete model privacy, supports resource-constrained mobile or edge devices, and significantly limits the attack surface for malicious clients to data poisoning only.
  • $\bullet$
    We systematically evaluate the performance of WW-FL instantiations as well as its resilience against state-of-the-art poisoning attacks. Additionally, we provide a prototype implementation and evaluation of WW-FL on standard image classification tasks.
In Tab. 1, we summarize how WW-FL distinguishes itself from related works.
2. Preliminaries and Related Work
This section provides a brief background on privacy-enhancing technologies such as MPC and HE, as well as an introduction to FL, along with a concise overview of related works.
Secure Multi-Party Computation (MPC). MPC (Yao, 1986; Goldreich et al., 1987) enables a set of mutually distrusting parties to evaluate a public function $f()$ on their private data while preserving input data privacy. The corruption among the parties is often modelled via an adversary that may try to take control of the corrupted parties and coordinate their actions. There are various orthogonal aspects of adversarial corruption like honest vs. dishonest majority, semi-honest vs. malicious corruption, etc. (Evans et al., 2018; Lindell, 2020). MPC is practically efficient when run among a small number of computing parties (Patra and Suresh, 2020; Knott et al., 2021; Keller and Sun, 2022; Koti et al., 2022). In our prototype implementation of WW-FL, we utilize Meta’s CrypTen framework (Knott et al., 2021), which for efficiency reasons operates in a semi-honest two-party setting with a trusted third “helper” party that generates correlated randomness (Riazi et al., 2018; Chaudhari et al., 2019; Patra et al., 2021).
Homomorphic Encryption (HE). HE schemes (Rivest et al., 1978; Gentry, 2009) enable computation on encrypted data without the need for decryption. Additively homomorphic encryption (AHE) is a widely used method that allows for the generation of a new ciphertext representing the sum of multiple plaintexts through operations on their corresponding ciphertexts (Paillier, 1999). In scenarios involving multiple parties, recent multi-party homomorphic encryption (MHE) schemes can reduce the communication complexity of MPC, but have high computation overhead (Mouchet et al., 2021). Furthermore, they are often secure only against semi-honest corruptions, whereas the recent work in (Chatel et al., 2023) is an exception. A comprehensive survey of various HE schemes is given in (Acar et al., 2018).
Differential Privacy (DP). The concept of DP (Dwork and Roth, 2014) is based on the idea of adding noise to data in order to reduce information leakage when sharing it, while still allowing for meaningful computations to be carried out on the data. Though DP techniques offer some protection against attacks in federated learning, they also reduce the utility of the data and the resulting ML model. Additionally, achieving robust and accurate models necessitates significant privacy budgets, leaving the actual level of privacy achieved in practice uncertain (Ouadrhiri and Abdelhadi, 2022). We refer to (Desfontaines and Pejó, 2020) for more details on DP.
Federated Learning (FL). Unlike conventional PPML techniques, FL (Konečnỳ et al., 2016; McMahan et al., 2017) enables the training of a machine learning model on distributed data by allowing each device to train the model locally using its own data. In each iteration, these local models are sent to a central server (also called an aggregator), where they are aggregated to form a global model. At a high level, an FL scheme iterates through the following steps:
  • $\bullet$
    The global server $\mathcal{S}$ sends the current global model $W_{t}$ to a selected subset of $n$ out of $N$ clients.
  • $\bullet$
    Each selected client $C^{i}$, $i\in[n]$ utilizes its own local training data $D^{i}$ for $E$ epochs to fine-tune the global model and obtains an updated local model $w_{t+1}^{i}$:
    $w_{t+1}^{i}\leftarrow W_{t}-\eta_{C^{i}}\frac{\partial L(W_{t},B_{i,e})}{\partial W_{t}},$
    where $L$ is a loss function, $\eta_{C^{i}}$ is the clients’ learning rate, and $B_{i,e}\subseteq D^{i}$ is a batch drawn from $D^{i}$ in epoch $e$, where $e\in[E]$. The local model updates $w_{t+1}^{i}$ are then sent back to $\mathcal{S}$.
  • $\bullet$
    $\mathcal{S}$ employs an aggregation rule $f_{\textit{agg}}$ to combine the received local model updates $w_{t+1}^{i}$, resulting in a global model $W_{t+1}$, which will serve as the starting point for the next iteration:
    $W_{t+1}\leftarrow W_{t}-\eta_{\mathsf{S}}\cdot f_{\textit{agg}}(w_{t+1}^{1},\ldots,w_{t+1}^{n}),$
    where $\eta_{\mathsf{S}}$ is the server’s learning rate.
The above procedure is repeated until a predefined stopping criterion is satisfied, e.g., a specified number of training iterations or a specific level of accuracy.
Secure Aggregation (SA). Standard FL’s privacy issues (Zhu et al., 2019; Geiping et al., 2020) led to single-server secure aggregation protocols (Single $\mathcal{S}$(Bonawitz et al., 2017; Mansouri et al., 2023) that use masking techniques or HE (Rubin, 1996; Bonawitz et al., 2017; Bell et al., 2020; Phong et al., 2018; Mandal and Gong, 2019). However, complete privacy for the global model remains elusive, posing risks if clients and aggregators collude. Moreover, recent research has shown that a single malicious server can reconstruct individual training data points from users’ local models even when using secure aggregation protocols (Pasquini et al., 2022; Boenisch et al., 2023; Zhao et al., 2023). Such vulnerabilities against colluding parties are addressed by multi-server SA (Multi $\mathcal{S}$(Fereidooni et al., 2021; Dong et al., 2021; Nguyen et al., 2022) and techniques like Spindle and Poseidon (Froelicher et al., 2021; Sav et al., 2021), albeit with scalability challenges. Recent works like (Dong et al., 2021) and (Nguyen et al., 2022) combine HE and MPC techniques to achieve private and robust FL, albeit at the cost of considerable computation and communication overhead in cross-device scenarios comprising hundreds to thousands of clients (Shejwalkar et al., 2022). In WW-FL, we employ a multi-server aggregation scheme, and mitigate scalability issues of existing works, especially for a heterogeneous cross-device setting.
Hierarchical Federated Learning (HFL). Another approach for FL, called hierarchical FL (HFL) addresses scalability and heterogeneity issues in real-world systems (Bonawitz et al., 2019; Deng et al., 2021; Wang et al., 2021). In HFL, aggregation occurs at multiple levels, forming a hierarchical structure in which the aggregated values from one level serve as inputs for the above level. This procedure eventually reaches the top level, where the final model is aggregated (Deng et al., 2021; Wang et al., 2021). Previous works in the HFL setting have primarily focused on improving scalability and communication. However, (Yang, 2021) introduced methods for ensuring the privacy of individual client updates in HFL, allowing for secure aggregation. Despite this advancement, (Yang, 2021) did not address global model privacy or robustness against malicious users, which we provide in WW-FL.
Poisoning & Robust Aggregation. One potential drawback of FL over PPML techniques is that FL increases the attack surface for malicious actors: in FL, each user trains their own model locally, allowing them to potentially manipulate their model (Biggio et al., 2012; Bhagoji et al., 2019). To counter these poisoning attacks, various robust aggregation methods have been proposed, in which local models are analyzed and appropriate measures such as outlier removal or norm scaling are applied (Yin et al., 2018; Li et al., 2019; Cao et al., 2021). See (Shejwalkar et al., 2022) for a comprehensive overview of different robust aggregations.
3. The WW-FL Framework
We now present the details of our WW-FL framework, which addresses multiple key requirements in FL: global model privacy, scalability, support for resource-constrained (mobile/edge) devices, reduction of attack surface, ability to defend against data poisoning, and high levels of user engagement. Our framework captures various proposed architectures for FL in a unified abstraction. We illustrate our framework in Fig. 2 and detail the underlying training algorithm in Alg. 1. The notations used in WW-FL are listed in Tab. 2.
[…]
$\mathsf{GS}_{i}\in\mathcal{G};i\in[\mathsf{N}_{\mathcal{G}}]$
Layer II: MPC Clusters
$\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i};i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
Layer III: Clients
$\mathsf{C}_{i}^{j}\in\mathcal{C}_{i};i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Training Data Adversary $\mathcal{A}$ MPC Server (Cluster) Private ML Training Cluster Models MPC Server (Global) Secure Aggregation Aggregated ML Model
Figure 2. Three-layer architecture in WW-FL for federated training of a machine learning model.
Table 2. Notations used in our WW-FL framework.
Notation Description
$\mathcal{G}$ Set of Global MPC Servers.
$\mathcal{M}_{i}$ Set of MPC Cluster Servers in the $i$-th cluster.
$\mathcal{C}_{i}^{\sf U}$ Set of Clients in the $i$-th cluster.
$\mathcal{C}_{i}$ Selected Clients in the $i$-th cluster.
$\mathsf{N}_{s}$ Size of the set $s\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}_{j}\}$.
$\mathsf{N}_{\footnotesize\mathcal{M}}$ Total number of clusters.
$W_{t}$ Global model available at round $t$.
$\mathsf{GS}_{i}$ Layer I MPC Global Server; $\mathsf{GS}_{i}\in\mathcal{G}$.
$\mathsf{CS}_{i}^{j}$ Layer II MPC Cluster Server; $\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i}$. Here, $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
$\mathsf{C}_{i}^{j}$ Layer III Client; $\mathsf{C}_{i}^{j}\in\mathcal{C}_{i}$, $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
${\langle\cdot\rangle}_{s}$ Secret sharing semantics for $s\in\{\mathcal{G},\mathcal{M}_{j}\}$.
3.1. WW-FL Architecture
Our WW-FL framework is based on a three-layer [^3] Additional layers can be easily incorporated depending on the scale of deployment. architecture and extends the established hierarchical FL (HFL) paradigm (Bonawitz et al., 2019; Wang et al., 2021; Yang, 2021). In HFL, clients are initially organized into clusters, and their local models are aggregated at the cluster level; these cluster-level models are then further aggregated globally, resulting in an additional layer of aggregation. We adopt the hierarchical approach in WW-FL for two main reasons: 1) to efficiently realize our distinct model training approach for global model privacy, and 2) to facilitate large-scale deployments across heterogeneous clients. However, there is one key difference: at the cluster level, WW-FL does not perform aggregation of models but private training on collective client data. This differs from both traditional and hierarchical FL since the data is not kept solely on the client devices. Nevertheless, WW-FL ensures that the data remains within the trusted domain of the client, allowing for accurate modeling of trust relationships between clients in real-world situations, such as trust among members of the same region or country (Marti and Garcia-Molina, 2006). Furthermore, this type of hierarchy is ubiquitous in real-world scenarios such as P2P gaming, organizations, and network infrastructures (Subramanian et al., 2002; Hong and Varghese, 2019; Lin et al., 2020).
We provide details for each layer in WW-FL next. We focus on the sequence of operations performed by the entities in our architecture (cf. Fig. 2) for a training over $T$ iterations while considering necessary MPC protocols and setup requirements. To make our design generic, we use the $\mathcal{F}_{\textsf{MPC}}$-hybrid model, a standard way to abstract MPC protocols. Concrete instantiations are given in Tab. 3 in §3.2.
Layer III: Clients. This layer is composed of $\mathsf{N}_{\footnotesize\mathcal{M}}$ distinct sets of clients $\mathcal{C}_{i}^{\sf U}$ (with $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]$), called clusters, which are formed based on specific criteria of the application (e.g., European Union (EU) member states for the EU smart metering scheme (Cuijpers and Koops, 2013; Commission, 2014)). Like standard FL, only a random subset of clients $\mathcal{C}_{i}\subseteq\mathcal{C}_{i}^{\sf U}$ will be selected by the training algorithm in an iteration $t\in[1,T]$.
During iteration $t$, client $\mathsf{C}_{i}^{j}\in\mathcal{C}_{i}$ (with $j\in[\mathsf{N}_{\mathcal{C}_{i}}]$) holding data ${D}_{t}^{\mathsf{C}_{i}^{j}}$ uses the Share protocol (cf. Tab. 3) to securely distribute its data to a set of MPC cluster servers $\mathcal{M}_{i}$. $\mathcal{M}_{i}$ are a representative group of high-performance servers that are trusted by their clients not to collude entirely, though some level of collusion among a subset of servers may occur. WW-FL allows clients to share their input and then leave at any time. They can also rejoin the system later and provide additional data in the next iteration they get selected. Hence, the clusters $\mathcal{C}_{i}^{\sf U}$ are dynamic and change with each iteration.
Our method differs from the standard concept of “data residing at the clients” in FL, but we expect it to not negatively impact user engagement as data remains within the users’ trust zone. Additionally, the reduced computational load allows resource-constrained devices to train complex models and eliminates the need for a shared-key setup among clients, simplifying dropout handling.
Layer II: MPC Clusters. The second layer consists of $\mathsf{N}_{\footnotesize\mathcal{M}}$ sets of distributed training servers $\mathcal{M}_{i}$ (with $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$), called MPC clusters, with each $\mathcal{M}_{i}$ corresponding to the cluster $\mathcal{C}_{i}^{\sf U}$ in Layer III. In iteration $t$, Layer I servers (denoted by $\mathcal{G}$) initiate ML training by sharing the current global model $W_{t-1}$ among the servers in $\mathcal{M}_{i}$. As will be discussed in §3.1, $W_{t-1}$ is also in a secret-shared form among $\mathcal{G}$, represented by ${\langle W_{t-1}\rangle}_{\mathcal{G}}$. To account for varying availability and trustworthiness of servers across regions, MPC clusters in WW-FL may use different MPC configurations and differ, e.g., in their corruption threshold and security model (Evans et al., 2018). Therefore, $\mathcal{G}$ uses the Reshare protocol (cf. Tab. 3) to convert the secret shares of ${\langle W_{t-1}\rangle}_{\mathcal{G}}$ to those of $\mathcal{M}_{i}$, i.e., ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$.
Given ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$, the servers in $\mathcal{M}_{i}$ use the Train protocol (cf. Tab. 3) to run MPC-based PPML for private ML training (Knott et al., 2021; Keller and Sun, 2022) on the cumulative secret-shared data from all clients in the cluster $\mathcal{C}_{i}$, denoted by ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}$. This data may include data from the same cluster in the previous iteration. Furthermore, by utilizing a larger pool of training data, we can leverage the known benefits of batching, resulting in faster convergence (Goyal et al., 2017; Bottou et al., 2018). After completing training, the servers in $\mathcal{M}_{i}$ run Reshare to secret-share the updated model with the Layer I servers, i.e., ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}$.
To preserve the system’s integrity, the servers for each MPC cluster must be chosen with care to ensure that clients are willing to share their data among the servers and that not all of the servers are colluding. One possible option is to build non-profit partnerships, such as in the MOC alliance (Zink et al., 2021), where organizations with mutual distrust can securely co-locate servers in the same data center with high-speed network connections. Alternatively, trusted entities like government organizations with limited infrastructure can host their servers in confidential cloud computing environments (Russinovich et al., 2021). If a client lacks trust in the WW-FL servers in its current cluster, the client can propose a new cluster with a partner that provides a non-colluding server.
Layer I: Global Servers. The top layer consists of a set of MPC servers $\mathcal{G}$, named Global Servers, that securely aggregate trained models from all the MPC clusters in Layer II, similarly to a standard FL scheme with a distributed aggregator (Fereidooni et al., 2021; Ben-Itzhak et al., 2022; Rathee et al., 2023; Gehlhar et al., 2023). Given the locally trained models $W_{t}^{i}$ for $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]$, the servers in $\mathcal{G}$ execute the secure aggregation protocol Agg  (Mansouri et al., 2023) to compute the updated global model in secret-shared form, i.e., ${\langle W_{t}\rangle}_{\mathcal{G}}$. The global servers in $\mathcal{G}$ use the Reshare protocol to distribute the aggregated model $W_{t}$ to each of the Layer II clusters $\mathcal{M}_{i}$ for the next iteration $t+1$.
The distributed aggregator model, as outlined in §1, circumvents the possibility of several privacy attacks that were recently presented for single-server aggregators even when relying on maliciously secure aggregation protocols (So et al., 2021; Boenisch et al., 2021, 2022; Fowl et al., 2022; Wen et al., 2022; Zhao et al., 2023); for example, by excluding certain inputs from the computation and thus influencing the correctness (but not the security) of the protocol, a malicious single server can mount privacy attacks that allow to infer information about particular participants and their training data. In contrast, in our distributed aggregator setting, even when having only semi-honest protocol security, such attacks are detectable as long as at least one of the servers behaves honestly.
Algorithm 1 FL-Train: Training in WW-FL
0:  $\mathcal{G},\mathcal{M},\mathcal{C}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$, $\mathcal{C}=\bigcup_{i}\mathcal{C}_{i}$; $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$
0:  $W_{0}$, $\{{D}^{\mathsf{C}}\}_{\mathsf{C}\in\mathcal{C}}$ $W_{0}$ – initial model, ${D}^{\mathsf{C}}$ – client $\mathsf{C}$’s data
0:  ${\langle W_{T}\rangle}$ $W_{T}$ – global model after $T$ iterations
1:  initialize: ${\langle W_{0}\rangle}_{\mathcal{G}}\leftarrow\textsc{Share}(W_{0},\mathcal{G})$
2:  for each training iteration $t\in[1,T]$ do
3:     for all $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$ do # in parallel
4:        ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}\leftarrow\mathcal{G}.\textsc{Reshare}({\langle W_{t-1}\rangle}_{\mathcal{G}},\mathcal{M}_{i})$
5:        $\mathcal{C}_{i}\leftarrow\mathcal{M}_{i}.\textsc{Sample}(\mathcal{C}_{i}^{\sf U},t)$$\mathcal{C}_{i}^{\sf U}$ – total clients in $i$-th cluster
6:        for all $j\in[\mathsf{N}_{\mathcal{C}_{i}}]$ do # in parallel
7:           ${\langle{D}_{t}^{\mathsf{C}_{i}^{j}}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathsf{C}_{i}^{j}}.\textsc{Share}({D}_{t}^{\mathsf{C}_{i}^{j}},\mathcal{M}_{i})$
8:        end for
[…]
10:        ${\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathcal{M}_{i}}.\textsc{Train}({\langle W_{t-1}\rangle}_{\mathcal{M}_{i}},{\langle{D}_{t}\rangle}_{\mathcal{M}_{i}})$
11:        ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{M}_{i}}.\textsc{Reshare}({\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}},\mathcal{G})$
12:     end for
13:     ${\langle W_{t}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{G}}.\textsc{Agg}(\{{\langle W_{t}^{i}\rangle}_{\mathcal{G}}\}_{i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]})$
14:  end for
WW-FL – The Complete Picture. Alg. 1 (FL-Train) provides the training algorithm in WW-FL. We detail the necessary MPC functionalities along with the WW-FL training functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ (Alg. 2) in §3.3. Although WW-FL has a 3-layer architecture, it can easily accommodate more levels of hierarchy depending on the size of the deployment. For this, additional layers of MPC clusters can be added between layers I and II, with the clusters performing secure aggregation instead of PPML training.
Existing schemes for global model privacy, such as (Mandal and Gong, 2019) and (Fereidooni et al., 2021), only protect the model from either clients or aggregator servers, leaving the possibility of collusion especially in a cross-device setting. WW-FL addresses this issue by keeping the global model in secret-shared form, ensuring that no single entity or group of colluding entities (up to an allowed corruption threshold) can access the model. This provides a stronger notion of privacy and also protects against unauthorized use or misuse, such as a client disclosing the trained model to another organization for further training or commercial use. For scenarios demanding disclosure of the global model to authorized entities, such as consortium members, the global servers can utilize the Reveal protocol.
3.2. MPC Functionalities in WW-FL
The MPC functionalities utilized in WW-FL are listed in Tab. 3 and discussed in the following. While Share and Reshare are used for generating the secret shares as per the underlying MPC semantics, Reveal is used to reconstruct the secret towards a designated party. The Train and Predict functionalities correspond to PPML training and inference protocols, respectively. Similarly, Agg denotes the secure aggregation functionality in FL (cf. §2). In WW-FL, these functionalities are realized using Meta’s CrypTen framework, in which two semi-honest MPC servers carry out the computation with the help of a trusted third party that provides correlated randomness (Damgård et al., 2012; Demmler et al., 2015; Riazi et al., 2018; Knott et al., 2021). However, WW-FL is not bound to any specific MPC setting and each of the MPC clusters as well as the global servers could be instantiated independently using any MPC protocol, as detailed further in §3.3 and §3.5.
Table 3. MPC functionalities used in WW-FL framework.
Algorithm Input(s) Description
Share ${D},{V}$ Generates secret shares of ${D}$ as per target ${V}$’s sharing semantics (${\langle{D}\rangle}_{{V}}$).
Reshare ${\langle{D}\rangle}_{{U}},{V}$ Converts secret shares of ${\langle{D}\rangle}$ from source ${U}$’s sharing semantics to target ${V}$.
Agg $\{{\langle{W}^{j}\rangle}\}$ Performs secure aggregation over $j$ secret-shared ML models.
Train ${\langle{W}\rangle},{\langle{D}\rangle}$ Performs PPML Training on ML model ${W}$ using the secret-shared data ${\langle{D}\rangle}$.
Predict ${\langle{W}\rangle},{\langle{Q}\rangle}$ Performs PPML Inference on secret-shared ML model ${\langle{W}\rangle}$ using the secret-shared query ${\langle{Q}\rangle}$.
Reveal ${\langle{D}\rangle}_{{U}},{V}$ Reconstructs secret-shared data ${\langle{D}\rangle}$ towards members in target set ${V}$.
TM-List $\mathcal{W},{\alpha}$ Performs Trimmed Mean and returns $2{\alpha}$ outlier values (top and bottom ${\alpha}$) for each index position in elements of $\mathcal{W}$.
TopK-Hitter $\mathcal{U},\Gamma$ Returns list of $\Gamma$ values that occur most frequently in $\mathcal{U}$.
3.3. Threat Model and Privacy Analysis
Given the hybrid nature of our framework, we consider a mixed adversarial corruption in WW-FL and assume a centralized adversary $\mathcal{A}$ orchestrating the corruption [^4] This is a stronger notion than having multiple independent adversaries., which allows the nature of the corruption to differ amongst MPC clusters including the global servers. We model this by defining a corruption strategy $\mathsf{St}_{\gamma}$ for the entities in the set $\gamma\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$. For example, $\mathcal{A}$ can choose to maliciously corrupt certain MPC clusters $\mathcal{M}_{i}$ (reflecting an unjustified trust assumption in the cluster), while corrupting the rest, including $\mathcal{G}$, in a semi-honest manner. $\mathcal{A}$ also corrupts a minority or majority of entities in a set, depending on the number of trustworthy entities (assuming at least one trustworthy entity in each set). The security spectrum of WW-FL is summarized in Tab. 4, ranging from weaker semi-honest cases to the strongest all-malicious setting based on the nature of entity corruption.
Table 4. Spectrum of security achievable in  WW-FL based on the corruption strategy of the adversary. Here $0<t<\mathsf{N}_{\mathcal{M}}$.
Security Global: $\mathcal{G}$ Clusters: $\mathcal{M}$ Clients: $\mathcal{C}$
I-A $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{semi\mbox{-}honest}$
I-B $\mathsf{semi\mbox{-}honest}$ $t$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
I-C $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
II-A $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{semi\mbox{-}honest}$
II-B $\mathsf{malicious}$ $t$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
II-C $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
III-A $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{malicious}$
III-B $\mathsf{semi\mbox{-}honest}$ $t$ $\mathsf{malicious}$ $\mathsf{malicious}$
III-C $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{malicious}$
IV-A $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{malicious}$
IV-B $\mathsf{malicious}$ $t$ $\mathsf{malicious}$ $\mathsf{malicious}$
IV-C $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{malicious}$
$\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ Ideal Functionality
We model the FL training process in WW-FL as an ideal functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ which is depicted in Alg. 2. For simplicity, we assume that the clients have secret-shared their training data with their respective MPC clusters via the respective Share protocol. In each iteration $t$, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives the shares of the current model $W_{t-1}$ from the global servers ($\mathcal{G}$) and the shares of the training data from each cluster’s servers in $\mathcal{M}_{i}$ for $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$. Using these shares, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ reconstructs the current global model and the training data for each cluster. Next, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ trains the model $W_{t-1}$ using the data from each cluster, resulting in a cluster model. These cluster models are then combined using an aggregation algorithm to obtain the updated global model $W_{t}$ for this iteration. Finally, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates the secret-shares of $W_{t}$ among the global servers ($\mathcal{G}$) and the cluster servers ($\mathcal{M}_{i}$) using their respective sharing semantics.
Algorithm 2 $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$: Ideal Functionality for WW-FL
0:  $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$, $\mathcal{G},\mathcal{M}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$; $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$
0:  ${\langle W_{t\mbox{-}1}\rangle}$, $\{{\langle{D}_{t}^{\mathcal{M}_{i}}\rangle}\}_{\mathcal{M}_{i}\in\mathcal{M}}$ $W_{t\mbox{-}1}$ – current global model, ${D}_{t}^{\mathcal{M}_{i}}$ – data at $\mathcal{M}_{i}$
[…]
1:  initialize: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives ${\langle W_{t\mbox{-}1}\rangle}$ from $\mathcal{G}$ and reconstructs $W_{t\mbox{-}1}$.
2:  for each MPC cluster $\mathcal{M}_{i}\in\mathcal{M}$ do # in parallel
3:     data reconstruction: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives ${\langle{D}_{t}^{\mathcal{M}_{i}}\rangle}$ from $\mathcal{M}_{i}$ and reconstructs the cluster data ${D}_{t}^{\mathcal{M}_{i}}$.${D}_{t}^{\mathcal{M}_{i}}$ corresponds to collective data at $\mathcal{M}_{i}$
4:     training: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ trains the current global model $W_{t\mbox{-}1}$ on ${D}_{t}^{\mathcal{M}_{i}}$ to obtain the cluster model $W_{t}^{\mathcal{M}_{i}}$.
5:  end for
6:  aggregation: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ aggregates $\{W_{t}^{\mathcal{M}_{i}}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$ using the respective secure aggregation algorithm to obtain $W_{t}{}$.
7:  secret-sharing (global servers): $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates ${\langle W_{t}\rangle}_{\mathcal{G}}$ from $W_{t}{}$.
8:  secret-sharing (cluster servers): $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates ${\langle W_{t}\rangle}_{\mathcal{M}_{i}}$ from $W_{t}{}$ for $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$.
The security of WW-FL is stated in Theorem 3.1, and we give detailed security proof in §A.2.
Theorem 3.1.
Algorithm FL-Train (Alg. 1) securely realises the $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ functionality (Alg. 2) in the $\mathcal{F}_{\textsf{MPC}}$-hybrid model, assuming the presence of a static adversary $\mathcal{A}$, who corrupts the entities in $\gamma\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$ using a predefined corruption strategy $\mathsf{St}_{\gamma}$.
3.4. Private Inference in WW-FL
In WW-FL, after the defined number of training iterations $T$ is completed, the MPC clusters serve as clusters for ML inference. Here, we again utilize PPML techniques to enable clients to privately query their clusters (Chaudhari et al., 2019, 2020; Knott et al., 2021; Mann et al., 2023). Consider the scenario where client $\mathsf{C}$ holding query ${Q}$ wants to use the inference service on a model $W$ that is secret-shared with a cluster $\mathcal{M}_{k}$. This is accomplished by $\mathsf{C}$ generating ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ using Share, followed by cluster servers in $\mathcal{M}_{k}$ running Predict on ${\langle W\rangle}_{\mathcal{M}_{k}}$ and ${\langle{Q}\rangle}_{\mathcal{M}_{k}}$ to generate the inference result in secret-shared form. Finally, $\mathcal{M}_{k}$ reveals the result to $\mathsf{C}$ using the Reveal protocol.
3.5. Abstraction of Existing FL Schemes
The abstraction of WW-FL captures multiple existing FL frameworks, including standard FL with a single aggregator (McMahan et al., 2017), distributed aggregators (Fereidooni et al., 2021), and hierarchical FL schemes (Yang, 2021), as shown in Tab. 5. This abstraction not only simplifies comparisons but also facilitates advanced hybrid designs, including the integration of differential privacy.
Table 5. Abstraction of existing FL schemes (cf. Tab. 1) in WW-FL (cf. Fig. 2). $\mathcal{S}$ – aggregation server(s), S.Agg. – secure aggregation, and (S.)Agg. – optional secure aggregation. See Tab. 2 for other notations.
Scheme Layer I Layer II Layer III Remark
Aggregation (Single $\mathcal{S}$) $\mathsf{N}_{s}$ $\mathsf{N}_{\mathcal{G}}=1$ $\mathsf{N}_{\mathcal{M}_{i}}=1$ $\mathsf{N}_{\mathcal{C}_{i}}=1$ $\mathcal{M}_{i}=\mathcal{C}_{i}$
Role (S.)Agg. ML Training
Aggregation (Multi $\mathcal{S}$) $\mathsf{N}_{s}$ $\mathsf{N}_{\mathcal{G}}>1$ $\mathsf{N}_{\mathcal{M}_{i}}=1$ $\mathsf{N}_{\mathcal{C}_{i}}=1$ $\mathcal{M}_{i}=\mathcal{C}_{i}$
Role S.Agg. ML Training
Hierarchical FL $\mathsf{N}_{s}$ $\mathsf{N}_{\mathcal{G}}=1$ $\mathsf{N}_{\mathcal{M}_{i}}=1$ $\mathsf{N}_{\mathcal{C}_{i}}>1$ $\mathcal{M}_{i}\neq\mathcal{C}_{i}$
Role (S.)Agg. (S.)Agg. ML Training
WW-FL This Work $\mathsf{N}_{s}$ $\mathsf{N}_{\mathcal{G}}>1$ $\mathsf{N}_{\mathcal{M}_{i}}>1$ $\mathsf{N}_{\mathcal{C}_{i}}>1$ $\mathcal{M}_{i}\neq\mathcal{C}_{i}$
Role S.Agg. PPML Training Data Sharing
Standard FL with a single aggregator (Single $\mathcal{S}$(McMahan et al., 2017) is a variant of WW-FL, where each Layer III cluster $\mathcal{C}_{i}$ consists of only one client that also serves as the MPC cluster server $\mathcal{M}_{i}$ in Layer II. Thus, it is sufficient to conduct ML training without global model privacy concerns and then send the results to a single global server $\mathcal{G}$ in Layer I for aggregation. The case of distributed aggregators (Multi $\mathcal{S}$(Fereidooni et al., 2021) follows similarly, except that secure aggregation is performed at Layer I with multiple ($\mathsf{N}_{\mathcal{G}}>1$) global servers. Finally, existing hierarchical FL schemes (Yang, 2021) have a similar three-layer architecture as WW-FL, but use a single server at both the global and cluster level ($\mathsf{N}_{\mathcal{G}}=1$, $\mathsf{N}_{\mathcal{M}_{i}}=1$). While WW-FL employs PPML training at the cluster-server level, hierarchical FL uses secure aggregation. Additionally, clients in the hierarchical FL approach perform local model training, as opposed to only data sharing in WW-FL.
Moreover, advanced hybrid designs in WW-FL could incorporate data from privileged clients, such as trusted consortium members, without hiding the global model. The client can train the model locally and share the gradients with the aggregation servers.
Another hybrid could include a collaborative training model like SafeNet (Chaudhari et al., 2022), which resists poisoning attacks. SafeNet defines the global model as an ensemble of verified models, with the majority prediction determining the final output. This setup can replace an MPC cluster and its associated clients in WW-FL. The ensemble models can be aggregated or subjected to MPC-based clustering (Nguyen et al., 2022) to obtain a single model. The remaining process follows the WW-FL architecture, replacing operations in SafeNet on the plain model with corresponding MPC counterparts. Also, the training in the MPC clusters, which is resilient to poisoning attacks, can be categorized as attested training. The global servers can use a simple/mildly robust aggregation scheme to incorporate updates from these clusters, decreasing the likelihood of error propagation to higher layers.
4. Performance Evaluation
We systematically and empirically evaluate the performance of WW-FL in terms of accuracy as well as computation and communication overhead. We use standard image classification tasks and build a prototype implementation using the MPC framework CrypTen (Knott et al., 2021). Future research can extend our evaluation to stronger security models (cf. §3.3), as well as more sophisticated ML models and training tasks.
4.1. Implementation
We implement WW-FL based on the CrypTen MPC framework developed by Meta (Knott et al., 2021). CrypTen provides a PyTorch-style interface and implements MPC operations with GPU support. Specifically, CrypTen implements semi-honest arithmetic and Boolean secure two- and multi-party computation protocols that use a trusted third “helper” party to generate correlated randomness. CrypTen provides a “simulation” mode where the specified computation is performed on a single node in plaintext yet simulates all effects that computation in MPC would have on the results (e.g., due to limited fixed-point precision and truncation). We leverage this mode to efficiently evaluate WW-FL’s accuracy and later the impact of data-poisoning attacks. For our run-time and communication measurements, we benchmark the full MPC protocols. In all our experiments, the fixed-point precision in CrypTen is set to 22 decimal bits (the maximum value recommended by the developers).
We use CrypTen to implement (a) private training on Layer II and (b) distributed aggregation on Layer I. CrypTen out of the box supports private inference between Layer II and III. We extend CrypTen with an identity layer to enable model conversions and re-sharing. Additionally, we extend the implementation of convolutional layers to enable full GPU-accelerated training for such model architectures. Moreover, we provide the necessary code to orchestrate the various parties and components, thereby creating a unified simulation framework.
Setup. Plaintext FL and CrypTen-based WW-FL simulations are run on a single computing platform with two Intel Xeon Platinum 8168 CPUs, 1.5TB RAM, and 16 NVIDIA Tesla V100 GPUs. For realistic results in a distributed MPC deployment with two computational and one helper party, we use three Amazon AWS g3s.xlarge instances with 4 vCPUs, and 8GB GPU memory on an NVIDIA Tesla M60. These instances are located in the same AWS availability zone (due to high costs associated with routing traffic between different zones), yet we simulate intra- and inter-continental network connections by setting the bandwidth and latency to 1Gbps/100Mbps and 20ms/100ms, respectively.
Tasks. Following prior work such as (Ben-Itzhak et al., 2022), we evaluate WW-FL on two standard image classification tasks: recognizing (a) hand-written digits using LeNet trained on MNIST (LeCun et al., 1998), and (b) objects in one of 10 classes using ResNet9 trained on CIFAR10 (Krizhevsky, 2009). We simulate 1000 clients overall from which 100 are randomly sampled during each training iteration. In the FL setting, all sampled clients provide their computed update directly to the aggregation server. In the WW-FL setting, each of the 10 MPC clusters has 100 associated clients from which each samples 10 clients at random. Each client has 200 data points assigned to it at random with duplicates allowed between clients.
For plain FL, we use batch size 8 and learning rate 0.005, and train locally for 5 epochs before central aggregation. To allow for a fair comparison between FL and WW-FL, we use a correspondingly scaled batch size of 80 and a learning rate of 0.05 as in (Goyal et al., 2017). Each client and MPC cluster performs 5 local training epochs with either a batch size of 8 or 80, respectively, following the recommendation in (Goyal et al., 2017). Tab. 6 summarises all the (hyper) parameters used during the training.
Table 6. Training parameters used in FL and WW-FL.
Parameter FL WW-FL
# clients 1000
# clients selected per round 100
# MPC clusters - 10
# clients per MPC cluster - 100
# clients per MPC cluster per round - 10
size of client datasets 200
learn rate 0.005 0.05
# local epochs 5
batch size 8 80
4.2. Results
In this section, we answer the following research questions to empirically evaluate performance and discuss our results:
  • Q1:
    What is the accuracy difference between FL (McMahan et al., 2017) and WW-FL (in plaintext)?
  • Q2:
    What is the impact on accuracy for WW-FL when moving from plaintext to (simulated) MPC?
  • Q3:
    What are the run-time and communication overheads of (MPC-based) WW-FL compared to FL?
Q1 – FL vs WW-FL. In Fig. 3, we compare the validation accuracy of FL and WW-FL for image classification tasks for 500 rounds. Here, WW-FL converges significantly faster than regular FL, e.g., after 500 rounds of training ResNet9 on CIFAR10, WW-FL reaches 85.68% validation accuracy, whereas regular FL only reaches 65.95%. Similarly, for MNIST, WW-FL achieves 98.95% accuracy compared to 98.72% for regular FL. We attribute this to WW-FL pooling training data at the cluster level and thus being able to exploit the known benefits of batching (Goyal et al., 2017; Bottou et al., 2018). Further plots for up to 2000 epochs are given in Fig. 7 in §A.3.1.
Figure 3. Validation accuracy for FL and WW-FL for 500 iterations (top: LeNet/MNIST, bottom: ResNet9/CIFAR10).
Q2 – Accuracy Impact of MPC. In Fig. 4, we compare the plaintext validation accuracy (cf. Q1) to our CrypTen simulation to measure the impact of MPC (i.e., fixed-point representation with 22 bit decimal representation and truncation). While there is a slight difference in initial rounds, both implementations quickly converge to almost the same validation accuracy, with only a small difference on the order of 0.1%.
Figure 4. Validation accuracy for FL and WW-FL in plaintext and MPC (CrypTen simulation) for LeNet/MNIST training.
Q3 – MPC Overhead. Finally, we evaluate the overhead of MPC for secure training and aggregation. For this, we measure the run-times and communication for one iteration of LeNet/MNIST training (i.e., 5 local epochs) in AWS for one cluster (with 1Gbps bandwidth and 20ms latency) and one iteration of global aggregation (with 100Mbps bandwidth and 100ms latency). The training on the cluster level takes 315s and requires 5.28GB inter-server communication, which is multiple orders of magnitude more than local plaintext training in PyTorch (which only takes 0.07s). The aggregation over 10 cluster inputs is very efficient with 0.023s run-time and has no communication overhead since only linear operations are required, which are local operations in MPC.
Additional overhead that must be considered for clients is sharing data with the training cluster servers. In our setup, clients on expectation have to upload 3.31MB and 9.86MB in total for 500 rounds of training for MNIST and CIFAR10, respectively. Furthermore, we have to account for sharing the trained models from training clusters to the aggregation servers. Given the number of model parameters and CrypTen sharing semantics, each training cluster must transfer 0.49MB and 39.19MB per server for LeNet and ResNet9, respectively. This clearly shows that it is significantly more efficient for participants to upload their training data in secret-shared form compared to down- and uploading model parameters for each training round. In our evaluation setup, the training clusters and the aggregation layer use the same MPC configuration, hence no interactive re-sharing is necessary.
5. Attacks & Mitigations
In FL, malicious participants can degrade accuracy through data or model poisoning attacks (Tian et al., 2022). Because WW-FL models are not accessible to clients, these attackers are limited to manipulating the training data they provide, resulting in either targeted, backdoor, or untargeted attacks. Our work is focused on untargeted FL poisoning, as it is most relevant to real-world applications and more difficult to detect (Shejwalkar et al., 2022; Fang et al., 2020). We therefore propose how to systematically evaluate the effectiveness of state-of-the-art data-poisoning attacks (Xiao et al., 2012; Fang et al., 2020; Shejwalkar et al., 2022; Tolpegin et al., 2020) in the WW-FL setting and propose a new robust aggregation scheme as a possible mitigation.
5.1. Data-Poisoning Attacks in WW-FL
In data-poisoning attacks, malicious clients can perform arbitrary manipulations on the training data. State-of-the-art attacks are based on label flipping, where clients keep the legitimate training samples, yet exchange the associated labels according to different strategies (Fang et al., 2020; Tolpegin et al., 2020; Shejwalkar et al., 2022).
To evaluate WW-FL, we implemented four label-flipping attacks detailed below. We assume a single attacker controlling all malicious clients, coordinating their actions. Only one attack occurs at a time, and if multiple clients have the same data sample, their poisoned labels will be the same. The assumption is that each malicious client poisons all its data samples to maximize the attack’s impact. The data is poisoned once and the labels remain constant throughout the model’s training.
  • $\bullet$
[…]
  • $\bullet$
    Targeted Label Flipping (TLF (Tolpegin et al., 2020)): Simply flips all labels from a source class to a target class. In WW-FL, we always set the source class as 0 and the target class as 1.
  • $\bullet$
    Static Label Flipping (SLF (Fang et al., 2020; Shejwalkar et al., 2022)): Uses a fixed permutation to determine the new label for each poisoned sample, given by the equation $\textit{new}\_\textit{label}=\textit{num}\_\textit{classes}-\textit{old}\_\textit{label}-1$.
  • $\bullet$
    Dynamic Label Flipping (DLF (Shejwalkar et al., 2022)): Utilizing a surrogate model to flip labels for each sample, our implementation aggregates data from all malicious clients to train a model with the same architecture as the one used in WW-FL training. Once trained, this model is utilized for inference and the labels are assigned to the least probable output determined by the surrogate model. The term “dynamic” is used as the labels rely on the trained model, and by altering the training configuration, the poisoned labels will be modified accordingly. The exact training settings of the surrogate model are given in Tab. 7.
    Table 7. Training parameters for surrogate model in DLF.
    Parameter Epochs Batch Size Learning Rate Momentum Weight Decay
    Values 50 128 0.05 0.9 0.000,5
5.2. New Robust Aggregation Scheme
The most common FL aggregation scheme, FedAvg, simply computes a (weighted) average of all inputs (cf. §2). In contrast, robust aggregation schemes detect and exclude outliers, thus a suitable mitigation against data poisoning. From the schemes surveyed in (Shejwalkar et al., 2022; Gehlhar et al., 2023), we identify “FLTrust” (Cao et al., 2021) and “Trimmed Mean” (TM) (Yin et al., 2018) as the ones with most efficient MPC implementations.
FLTrust (Cao et al., 2021) measures the cosine similarity between the inputs of participants and the most recent global model trained with a clean data set, then excludes the least similar inputs. Trimmed Mean (TM) (Yin et al., 2018), for each coordinate, computes the mean across the provided gradient updates and excludes the values that deviate the most in either direction of the mean. Performing TM aggregation obliviously in MPC requires implementing costly sorting to determine the ranking in each coordinate.
We observe that, intuitively, data poisoning in contrast to model poisoning does not result in specific coordinates producing extreme outliers. Hence, we propose a heuristic “Trimmed Mean Variant” that computes the mean and ranking only for a small randomly sampled subset of coordinates. Then, during aggregation, it excludes those inputs that occurred the most as outliers in the sample. We detail our algorithm in Alg. 3, where the underlying MPC functionalities TM-List and TopK-Hitter are as follows:
Algorithm 3 Our Trimmed Mean (TM) Variant in WW-FL
0:  $\mathcal{W}=\{W^{i}\}_{i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]}$, ${\alpha}$, ${\beta}$, ${\gamma}$ ${\gamma}=|W^{i}|$, ${\alpha}$ – trim threshold, ${\beta}$ – sample size
0:  ${\langle W_{AGG}\rangle}$ aggregated model after removing outliers
1:  initialize: $\mathcal{Z}\leftarrow\emptyset$ set of outliers
2:  $\mathcal{I}\leftarrow$ sample random ${\beta}$ indices from $[1,{\gamma}]$.
3:  $\mathcal{U}\leftarrow\textsc{TM-List}(\mathcal{W}^{\mathcal{I}},{\alpha})$ $\mathcal{W}^{\mathcal{I}}$–Truncated $\mathcal{W}$ with only indices in $\mathcal{I}$, TM-List performs Trimmed Mean algorithm and returns $2{\alpha}$ outlier values (top and bottom ${\alpha}$) for each index in $\mathcal{I}$, $|\mathcal{U}|=2{\alpha}{\beta}$
4:  $\mathcal{V}\leftarrow\textsc{TopK-Hitter}(\mathcal{U},2{\alpha})$ returns list of $2{\alpha}$ indices that occur most frequently in $\mathcal{U}$
5:  for all $i\in\mathcal{V}$ do
6:     $\mathcal{Z}\leftarrow\mathcal{Z}\bigcup\{W^{i}\}$
7:  end for
8:  ${\langle W_{AGG}\rangle}\leftarrow\textsc{Agg}(\mathcal{W}\setminus\mathcal{Z})$
TM-List takes as input a set of vectors, say $\mathcal{W}$, consisting of ${\beta}$-sized vectors of the form $W_{j}^{i}$ for $i\in[{\beta}],j\in[\mathsf{N}_{\mathcal{W}}]$. Moreover, the values in the vector come from a fixed source (the Layer II MPC clusters in our case) and are thus represented as a tuple of the form $W_{j}^{i}=(u_{i},v_{i})_{j}$. Here, $u_{i}$ denotes the source ID (MPC cluster in WW-FL) and $v_{i}$ represents the corresponding value. W.l.o.g., consider the first index position of these vectors ($i=1$). TM-List sorts the list $\{(u,v)_{j}\}_{j\in[\mathsf{N}_{\mathcal{W}}]}$ using the value $v$ as the key and selects the IDs ($u$) associated with the top and bottom ${\alpha}$ values. Intuitively, the operation results in selecting the MPC clusters whose local updates fall in either the top-${\alpha}$ or bottom-${\alpha}$ position among all the updates at that index. This procedure is performed in parallel for all ${\beta}$ indices and results in a set $\mathcal{U}$ of $2{\alpha}{\beta}$ IDs (with duplicates). The TopK-Hitter functionality, parameterized by $\Gamma$, takes this set $\mathcal{U}$ as input and returns a set of $\Gamma$ values that occur most frequently in $\mathcal{U}$.
Implementation Details
CrypTen lacks a built-in oblivious sorting functionality, so we implement privacy-preserving sorting for so-called CrypTensors. Sorting is necessary to compute trimmed mean and our optimized trimmed mean variant. We minimize the number of comparisons by implementing a bitonic sorting network that generates the necessary comparisons between elements to allow for a non-stable sorting algorithm. For trimmed mean, it is not necessary to preserve the relative order of same valued keys as each of them would have been seen as suspicious anyway. The comparisons are stored in plaintext, as they do not reveal any information about the data. The comparison steps are only computed once and then executed in parallel for each coordinate. For 100 elements, we perform 1077 comparisons and for 10 elements only 31. As the result of each comparison is hidden, we perform the swap operation for each pairs as described in Lst. LABEL:listing:crypten-sorting.
1compare_indices = list(comparision_generator(tensors.size(1)))
2for (i, j) in compare_indices:
3 b = tensors[:, i].gt(tensors[:, j])
4 h = b * tensors[:, i] + (1 - b) * tensors[:, j]
5 l = b * tensors[:, j] + (1 - b) * tensors[:, i]
6 tensors[:, i] = l
7 tensors[:, j] = h
Listing 1: CrypTen: Parallel Sorting.
When computing the proposed trimmed mean variant, we are only interested in the indices of the outliers and therefore only perform the swap operations on a list of indices. This is done to minimize the compute operations and thereby reduce time and communication. After identifying which gradients were most often detected as outliers and then computing the set of benign indices, we finally compute the sum over those while preserving privacy. The procedure is shown in Lst. LABEL:listing:sumofbenign.
1def bincount(self, indices, num_bins):
2 indices_list = crypten.cryptensor(torch.arange(0, num_bins))
3 counts = crypten.stack([indices for _ in range(num_bins)],dim=1).eq(indices_list).
4 sum(dim=0)
5 return counts[None, :]
6
7def exclude_topk_indices(self, indices, k):
8 sorted_indices = self.sort(indices).reshape(-1)
9 n = sorted_indices.shape[0]
10 return sorted_indices[:n-k]
11
12benign_indices = self.exclude_topk_indices(
13 self.bincount(outlier_indices, num_bins=n),
14 k=self.exclude_topk)
15
16all_indices = crypten.cryptensor(torch.arange(0, n).repeat(benign_indices.size(0), 1)).t()
17agg_sum = gradients.matmul(all_indices.eq(benign_indices)).
18 sum(dim=1)
19
20return agg_sum / (n - self.exclude_topk)
Listing 2: CrypTen: Sum of Benign Updates.
Tab. 8 shows the parameters used by the aggregation server. The trimmed mean threshold ${\alpha}$ has been chosen such that it can exclude all malicious updates on either side in the FL setting. The worst-case scenario we consider is a poison rate of 0.2, which is equivalent to approx. 20 malicious clients selected per round.
FLTrust (Cao et al., 2021) uses a root data set to calculate the server model. To be most effective, the server model must be similar to the benign models and therefore the root dataset must be representative for the whole dataset. Therefore, we sample (like for all clients) 200 data points for that dataset.
Table 8. Robust aggregation parameters.
Parameter FL WW-FL
Trimmed Mean ${\alpha}$ 20 2
Trimmed Mean Variant excluded gradients 40 4
FLTrust root data set size 200
Figure 5. Validation accuracy for FL and WW-FL when training ResNet9 on CIFAR10 with FedAvg, FLTrust, and trimmed mean as aggregation schemes under DLF attack for three different poison rates (top: equally distributed, bottom: focused setting).
5.3. Evaluation
We now empirically evaluate the impact of data-poisoning attacks on WW-FL considering different (robust) aggregation schemes. For this, we implement all four attacks—RLF (Xiao et al., 2012), TLF (Tolpegin et al., 2020), SLF (Fang et al., 2020; Shejwalkar et al., 2022), DLF (Shejwalkar et al., 2022)—in our framework and add CrypTen-based implementations of three robust aggregation schemes to our prototype implementation. Using the setup described in §4, we answer the following questions next:
  • Q4:
    What is the impact of data-poisoning attacks on the accuracy of FL and WW-FL using FedAvg and robust aggregation schemes?
  • Q5:
    What is the run-time and communication overhead for different robust aggregation schemes in MPC?
  • Q6:
    How does our TM variant compare to regular TM w.r.t. accuracy and MPC performance?
Q4 – Attack Impact. We consider three poison rates (0.01, 0.1, 0.2) and distributions of attackers: in the equally-distributed setting, we assume that on expectation each cluster has the same number of malicious clients; in the focused setting, we assume that malicious clients are concentrated on as few clusters as possible while there is still an honest majority in each cluster (a standard assumption in FL); finally, in the unrealistic cluster-focused setting, we see what happens if we lift the honest-majority assumption and concentrate all malicious clients in as few clusters as possible. In Fig. 5, we study the effectiveness of the most powerful DLF (Shejwalkar et al., 2022) attack on both regular FL and WW-FL when training ResNet9 on CIFAR10 in the equally distributed and focused setting. Additional experimental outcomes, including evaluation over the cluster-focused setting, are given in §A.3.2.
For the fairly aggressive poison rate of 0.2, we see in both visualized attacker distributions a significant negative impact of the DLF attack on FL when using FedAvg with drops below 30% accuracy. However, these can be successfully mitigated with robust aggregation schemes. While there is also negative impact on WW-FL, especially in the focused setting, the accuracy even with FedAvg never drops below that of FL. Even though robust aggregation schemes help to slightly smoothen the curve, we conclude that applying defenses in WW-FL against data-poisoning attacks is optional, but not strictly necessary.
Q5 – Robust Aggregation in MPC. We evaluate the run-time and communication overhead of our FLTrust and TM implementation in CrypTen in Tab. 9. The run-time overhead for both robust aggregation schemes compared to FedAvg is four to five orders of magnitude higher (ranging from 5 to 30 minutes). Also, FLTrust requires 5$\times$ more run-time and communication than TM. Given that both produce fairly similar results when applied to WW-FL, the overhead for FLTrust seems not justified.
Table 9. Communication (Comm. in MB) and run-time (Time in seconds) for various aggregation schemes, including our Trimmed Mean (TM) variant with sample sizes 10, 100, and 1000 in CrypTen.
Param FedAvg FLTrust TM TM-10 TM-100 TM-1000
Comm. 0 5,329.19 1,021.59 9.69 11.90 33.94
Time 0.023 1,713.44 326.57 558.05 558.26 561.50
Q6 – Trimmed Mean Variant. In Fig. 6, we compare the effectiveness of our TM variant to the original TM (Yin et al., 2018) for three sample sizes (10, 100, and 1000). It turns out that our heuristic approach barely reduces the effectiveness, even with aggressive parameters. In fact, in the focused setting, the TM variant outperforms the original. This is because our variant completely excludes updates of (poisoned) outliers, whereas in regular trimmed mean, those poisoned updates might still be considered for some coordinates. Results for all other settings are presented in §A.3.3.
Figure 6. Accuracy of trimmed mean (TM) and our variant (with sample sizes 10, 100, and 1000) against focused DLF attacks on WW-FL at 0.2 poison rate for ResNet9/CIFAR10.
In Tab. 9, we provide run-times and communication results for our optimizations. Compared to the original with 1.02GB of communication, we can see an improvement by two orders of magnitude for the variant with 100 random samples. However, we see a higher and fairly stable run-time across all three examined variants. This is because the algorithm for determining the overall ranking of outliers increases the number of MPC communication rounds. In the studied inter-continental WAN setting, this has severe impact but does not correspond to actual compute time. Overall, if WW-FL is combined with a robust aggregation scheme, our TM variant offers an excellent trade-off between accuracy and MPC overhead.
Additional Benchmarks. We give the detailed results of our evaluation in §A.3, including additional accuracy comparisons of WW-FL with the standard FL in all three modes of corruption, and the accuracy results for our Trimmed Mean variant.
6. Conclusion
In this work, we presented WW-FL, a novel unified abstraction and framework for large-scale (hierarchical) federated learning that provides global model privacy, faster convergence, smaller attack surface, and better resilience against poisoning attacks than regular FL. We envision that WW-FL will lead to a paradigm shift in global FL system design and deployment. Future efforts can extend our foundation work with efficient instantiations, stronger security, and comprehensive evaluations with larger datasets and setups.
Acknowledgements
This project received funding from the ERC under the EU’s Horizon 2020 research and innovation program (grant agreement No. 850990 PSOTI). It was co-funded by the DFG within SFB 1119 CROSSING/236615297 and GRK 2050 Privacy & Trust/251805230.
Appendix A Additional Details for WW-FL
This section provides additional details for our WW-FL framework, including motivation, security, and benchmarks.
A.1. Motivation for WW-FL
We elaborate on the motivation behind our WW-FL framework, which is specifically designed to facilitate the global deployment of FL. At a high level, WW-FL aims to address the diversity in trust levels among entities across different geographical locations (OECD, 2022; Statista Research Department, 2022). According to a survey (Statista Research Department, 2022), the global average trust in government was reported at $51\%$ in 2022. Furthermore, the degree of trust varies among different entities, such as law enforcement agencies and judicial systems (OECD, 2022).
When it comes to private companies, some countries have implemented strict laws to ensure the ethical behavior of entities operating within their borders. Non-compliance can result in substantial fines and even imprisonment (Husain, 2023; Hill, 2023). Recent examples include Meta receiving a 1.2 billion EUR fine from regulators in Ireland (Husain, 2023) and Didi Global facing a 1.19 billion EUR penalty from the Cyberspace Administration of China (Hill, 2023). On the other hand, several companies including well-known TalkTalk and Target either lost their customers’ data in major data breaches or were accused of not keeping customer information safe. This lead to a lack of trust among their clients and thereby face a huge drop in its revenue as well as loosing their customer base (Erin, 2022; Hibberd, 2022).
Given this disparity in trust, consider a consortium comprising three companies interested in training a machine learning (ML) model for medical diagnostics. Numerous studies have demonstrated the effectiveness of ML models in diagnosing various diseases, such as lung cancer, tumor detection, and liver segmentation (Rauniyar et al., 2023). However, these studies emphasize the need for diverse patient data. Simultaneously, the sensitivity and privacy of the training data must be preserved.
To address this, the consortium may collaborate with various federal, and local government agencies to obtain relevant training data and potentially compensating data providers. However, since the model is trained on sensitive medical data, the consortium members do not want the trained model to be disclosed to anyone. Additionally, the consortium should have the ability to use the trained model for inference while preserving privacy. This approach not only enhances the system’s privacy by safeguarding against breaches that might affect individual companies but also safeguards the financial investments of these companies by preventing unauthorized use of their trained model by third parties.
At first glance, implementing a Privacy-Preserving Machine Learning (PPML) solution may appear simplistic. In PPML, data owners securely distribute their data among the three companies within the consortium, ensuring that no individual company can access the data in the clear. The consortium employs MPC techniques to privately train a model on the shared data. However, this approach has several drawbacks. One significant drawback is the substantial computational power required within the consortium to handle the training process, particularly when dealing with millions of data owners in a global deployment. Furthermore, this approach necessitates a high level of trust from all data owners, as they must share their data among the consortium’s member companies. This trust requirement can be a significant hurdle to adoption. Additionally, there may be situations where data owners are willing to share their data but cannot do so due to government regulations preventing data from leaving their jurisdiction. This restriction not only limits data availability but also renders this approach unsuitable for our specific use case.
The next approach we can consider is using a traditional FL with secure aggregation. While this method has the potential to address the scalability problem, it comes with a significant drawback. In this approach, the trained model becomes publicly accessible to the data owners, which does not align with our specific scenario. Although there are works like (Froelicher et al., 2021; Sav et al., 2021), which protect the privacy of the global model using homomorphic encryption techniques, they may not be well-suited for our cross-device setting with a large number of data owners. Additionally, these techniques necessitate interaction among the clients and require extra measures for handling dropouts, as indicated in Tab. 1.
In light of the limitations of current methods, we have made a slight adjustment to the concept of FL in WW-FL. Specifically, we have expanded the notion of a trust zone, allowing data to be moved from the data owner to this larger, still-trusted area. For instance, a user in our scenario can securely share their data among three entities: the police, the court, and a private company, all operating within the same legal jurisdiction as the data owner. These selected entities will collaboratively train the model using privacy-enhancing technologies like MPC and HE. Furthermore, this approach has the added benefit of supporting edge nodes with limited computational resources, allowing them to contribute to the system. Importantly, since data owners can be assured that their data remains within their designated trust zone, this will likely encourage more users to participate, leading to a larger pool of training data and, ultimately, a better-trained model with improved global privacy.
Finally, it is important to understand that the concept of a trust-zone can differ depending on the specific region or context. For example, in one country, it may be relatively straightforward to identify three entities that data owners trust not to collude with each other. In contrast, there may be situations where it is challenging to find three or four entities that can provide an absolute guarantee of non-collusion. However, in such cases, data owners may be willing to accept a weaker assurance that not all entities will collude. WW-FL can accommodate this diversity in trust by performing PPML training independently for each of these scenarios, using the appropriate MPC protocol, and then combining their results through secure aggregation using global servers.
A.2. Security of WW-FL
This section provides security details of our WW-FL architecture from §3.1.
See 3.1
Proof Sketch
In the FL-Train algorithm, the initial step involves generating secret shares of the global model and distributing them among the global servers ($\mathcal{G}$). This step ensures correctness and privacy through the Share algorithm, a part of the $\mathcal{F}_{\textsf{MPC}}$ functionality. Once the global model is available in this secret-shared format, the next step is to securely redistribute it among servers within each cluster.
For example, consider the case where there are only two semi-honest global servers among which the global model is shared upon. Then, for an MPC cluster consisting of three servers (3PC), designed to tolerate one malicious corruption, the global servers employ the Reshare protocol in $\mathcal{F}_{\textsf{MPC}}$. This protocol converts their secret shares into the semantics suitable for the 3PC malicious cluster. Here again, the security relies upon the guarantees of the $\mathcal{F}_{\textsf{MPC}}$ ideal functionality.
The next step in FL-Train is for the clients to securely share their data with the cluster servers in their respective MPC clusters. This secure sharing is achieved through the Share protocol, which guarantees that the cluster servers will possess consistent shares of the data. Even in the event of malicious corruption, Share ensures that a malicious client cannot introduce inconsistent shares among the cluster servers. Note that a malicious client can tamper with its data before sharing it with an MPC cluster. However, this behavior is allowed in the $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ functionality, as $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ guarantees the privacy and correctness of the aggregation operation of the secret-shared data by the clients. The underlying secure aggregation algorithms (Agg) handle and address such attacks, which fall outside the scope of MPC. Nonetheless, $\mathcal{F}_{\textsf{MPC}}$ ensures the secure evaluation of these aggregation algorithms.
After the data and the current global model are distributed in secret-shared form among the cluster servers, they execute the Train algorithm to securely acquire the secret-shares of the trained cluster model. Subsequently, these cluster models are shared secretly among the global servers using the Reshare protocol. It is important to highlight that the security of these two steps is ensured by the $\mathcal{F}_{\textsf{MPC}}$ functionality. In the event of malicious corruption within the cluster, MPC protocols capable of tolerating such corruption will be employed. This will effectively detect any cheating attempt by a malicious cluster server with a very high probability.
After the cluster models have been transformed into secret shares, the global servers ($\mathcal{G}$) use the Agg algorithm to securely aggregate them and obtain the updated global model in secret-shared form. Note that the Agg algorithm, as mentioned before, plays a crucial role in defending against poisoning attacks from malicious clients and is outside the scope of MPC. However, it ensures the privacy and accuracy of the aggregation process, resulting in the correct global model. Finally, the global servers uses Reshare protocol to re-share the global model back to the cluster servers in their respective sharing semantics.
In summary, the correctness of Share, Reshare, Train, and Agg algorithms relies on the $\mathcal{F}_{\textsf{MPC}}$ functionality when instantiated correctly for the desired level of security. Therefore, with a very high probability, a malicious server within a cluster or at the global level cannot introduce errors into the system without being detected. At a high level, the underlying MPC protocols employ techniques such as error correction or replication in honest-majority settings, while dishonest-majority scenarios utilize message authentication codes to detect such errors. $\square$
A.3. Additional Benchmarks
In this section, we present additional benchmarks and results to further evaluate the performance of WW-FL. These supplementary findings provide valuable insights into the effectiveness of WW-FL in comparison to regular federated learning (FL) across various scenarios and attack settings.
A.3.1. Additional Results for Q1
Fig. 7 shows validation accuracy of FL and WW-FL for image classification tasks over the full 2000 training iterations.
Figure 7. Validation accuracy for FL and $WW\mbox{-}FL$ for 2000 epochs (top: LeNet/MNIST, bottom: ResNet9/CIFAR10)
A.3.2. Additional Results for Q4
ResNet9/CIFAR10
Here, we present the full evaluation results for ResNet9 trained on CIFAR10 under attack. We evaluate FedAvg, FLTrust, and trimmed mean for 2000 rounds in regular FL and WW-FL under four different data-poisoning attacks with three different poison rates and three different distributions of malicious clients. Fig. 8 shows the results for equally-distributed malicious clients, Fig. 9 the focused attack, and Fig. 10 cluster-focused results.
WW-FL outperforms regular FL in nearly all settings and has more than 5% higher validation accuracy after 2000 rounds. In both regular FL and WW-FL, all three aggregation schemes show robustness against data poisoning in the realistic settings with 0.01 and 0.1 poison rate and equally-distributed as well as focused attacks. This is not the case for a poison rate of 0.2. Especially FedAvg as a non-robust aggregation scheme struggles to converge. Even robust aggregations cannot fully counteract the attacks. The same is true for the cluster-focused attack distribution: Here, all aggregations have spikes in their validation accuracy; only FLTrust manages to train a model comparable to the other settings. This is because FLTrust is designed to be robust against model poisoning, and the cluster-focused setting is the closest to a model-poisoning attack. We stress that the cluster-focused scenario is less practical in reality. Consider worldwide training that is organized into country-level clusters: it is extremely unlikely that an adversary can corrupt even the majority of a country’s population. Corruption in a production-level cross-device setting is below a rate of 0.001 (Shejwalkar et al., 2022), making our study of extreme cases with a 0.2 corruption rate a theoretical experiment.
Figure 8. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 9. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 10. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 11. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 12. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 13. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 14. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 15. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 16. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
LeNet/MNIST
We show additional evaluation results for LeNet trained on MNIST for 2000 iterations in Fig. 11 (equally-distributed), Fig. 12 (focused), and Fig. 13 (cluster-focused). As for ResNet9, WW-FL trains the model faster than regular FL, but both converge to a similar accuracy in most cases after 2000 rounds. Here, the advantage of WW-FL over regular FL mostly lies in the faster convergence: after only 100 rounds of training, WW-FL nearly reaches 98% accuracy, regardless of the attack or attack setting. Also, there is little difference between the three aggregation schemes for the realistic settings. For the 0.2 poison rate, FedAvg starts to struggle a lot to train the model. This can mostly be seen for FedAvg in regular FL and sometimes in WW-FL.
A.3.3. Additional Results for Q6
ResNet9/CIFAR10
We present the extended evaluation of our trimmed mean variant for CIFAR10 when sampling 10, 100, and 1000 coordinates under all attacks and attack settings. Again, the plots show the evaluation results for 2000 iterations and are divided into three distributions for malicious clients: Fig. 14 shows equally-distributed, Fig. 15 shows focused, and Fig. 16 shows cluster-focused.
In most cases, all four aggregations perform equally and the final validation accuracy is nearly the same. The only difference can be seen with the cluster-focused attack distribution and 0.2 poison rate. Here, trimmed mean achieves a substantially lower accuracy than the trimmed mean variant. We assume that is because cluster-focused acts more like a model poisoning attack and trimmed mean variant deals better with this by discarding whole gradients rather than operating coordinate-wise.
LeNet/MNIST
We also compared trimmed mean against our trimmed mean variant for LeNet being trained on MNIST. However, almost all settings, the aggregations perform very similar. Hence, we omit the full plots.